mikewwe13:

Jason paid $15.50 for 3 slices of pizza and 2 burgers. Susan paid $20 for 1 slice of pizza and 4 burgers. Fill in the following:(Do NOT include a $ in your answer) Each slice of pizza costs $ ____________________. Each burger costs $ ______________________

mikewwe13:

@Vocaloid

mikewwe13:

@dude

dude:

We can set x = pizza and y = burgers For Jason we can say: 3x+2y=15.50 For Susan we can say: x+4y=20

mikewwe13:

i have a pic

dude:

Sure, make sure to include any image that relates to the question!

mikewwe13:

1 attachment
dude:

We can solve the 2 equations using the system of equations \(3x+2y=15.50\\ x+4y=20\) Do you want to solve by substitution of by elimination?

dude:

correction, or*

mikewwe13:

sure but do you know what those two last lines mean ?

mikewwe13:

Here's the hints Hint 1 Write the two equations that represent the given information. Let p = cost of a slice of pizza and let b = cost of a burger. Keep in mind how many pizzas and burgers each person bought. Hint 2 The equation for 3 pizzas and 2 burgers is 3p+2b=15.50. Hint 3 To solve, you will add the two equations. Use the distributive property to multiply one of the equations such that when you add them, one of the terms cancels out.

dude:

Sure, we did the same thing only with different variables \(3x+2y=15.50\\ x+4y=20\) We can multiply the second equation by -3 to get rid of x \(3x+2y=15.50\\ \color{red}{-3(}x+4y=20\color {red}{)}\)

mikewwe13:

those are the equations for the two lines ?

dude:

Yep

dude:

After distributing we get \(3x+2y=15.50\\ −3x-12y=-60\) now we just add both equations

mikewwe13:

ok from the beginning Each slice of pizza costs $ ____________________________ . Each burger costs $ _____________________

dude:

Do you want me to re-explain what I just did?

mikewwe13:

yes please not to be confusing

dude:

Alright so we have to get 2 equations that represent the cost of the burgers and pizza Jason and Susan ate

mikewwe13:

ok

dude:

We can set p= pizzas and b= burgers

mikewwe13:

k

dude:

Jason ate 3 slices of pizza and 2 burgers, so we can write 3x+2y We also know that the cost for everything he ate was $15.50 so we can set that equal to the number of items bought \(3x+2y=15.50\)

mikewwe13:

do i right this for jason 3x + 2y = 15.50 or just 15.50 ?

dude:

3x + 2y = 15.50 This one

mikewwe13:

ok now for burger

dude:

Hmm, y is the burgers Did you mean Susan?

mikewwe13:

yes sir

dude:

So we know that Susan ate 1 slice of pizza and 4 burgers, so we can write 1x+4y The cost for everything she bought was $20, so we set it equal to the items eaten \(\color{red}{1x+4y=20}\)

mikewwe13:

1x + 4y = 20 for burger ?

dude:

Susan* yes

mikewwe13:

ok for the two lines i write these right 3x + 2y = 15.50 1x + 4y = 20

dude:

Right

dude:

So now we solve using system of equations

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mikewwe13 The coefficients of the terms must be opposites, so they will have a sum of 0. \(\color{#0cbb34}{\text{End of Quote}}\)

mikewwe13:

ok that's it ?

mikewwe13:

thx a lot i appreicate it dude

dude:

No, we still need to solve for x and y

dude:

Ahhh that was not the answer!